You are operating a firm in a perfectly competitive market. In the short run, you have fixed costs of $30. Your variable costs are given in the following table:

Q TVC
0 0
1 100
2 150
3 180
4 220
5 300
6 390

Complete the following table:

Market Price Profit maximizing level of output Profit
$48
$60
$75
$85
Answer #1
Q TVC TFC TC MC
0 0 30 30
1 100 30 130 100
2 150 30 180 50
3 180 30 210 30
4 220 30 250 40
5 300 30 330 80
6 390 30 420 90

Setting P=MC, the output produced:-

Profit = TR-TC

Market Price Profit-maximizing level of output Profit
$48 4 48*4-250 = -58
$60 4 60*4-250 = -10
$75 4 75*4-250 = 50
$85 5 85*5-330 = 95